OpenStudy (anonymous):

Solve the differential equation (i.e find the equation for f(x)). 1.)f'(s)=10s-12s^3,f(3)=2 2.)f"(x)=sinx,f'(0)=1,f(0)=6

OpenStudy (anonymous):

\[\begin{align*}\frac{df(s)}{ds}&=10s-12s^3\\ \int df(s)&=\int(10s-12s^3)~ds\\ f(s)&=\frac{10s^2}{2}-\frac{12s^4}{4}+C\\ f(s)&=5s^2-3s^4+C \end{align*}\] Given that \(f(3)=2\), you have \[2=5(3)^2-3(3)^4+C\] Solve for \(C\).

OpenStudy (anonymous):

\[\begin{align*} f''(x)&=\sin x\\ \int f''(x)~dx&=\int\sin x~dx\\ f'(x)&=-\cos x+C_1\\ \int f'(x)~dx&=-\int(\cos x-C_1)~dx\\ f(x)&=-\sin x+C_1x+C_2 \end{align*}\] Given that \(f'(0)=1\) and \(f(0)=6\), you have \[\begin{cases} 1=-\cos 0+C_1\\ 6=-\sin 0+C_1(0)+C_2 \end{cases}\] Solve for \(C_1\) and \(C_2\).